OpenStudy (anonymous):

what is the value of x? x/4-8=12 x=?

OpenStudy (anonymous):

add 8 to both sides first.

OpenStudy (anonymous):

x/4- 8 +8=12 +8 x/4=20 now multiply both sides by 4 x/4*4=20*4 x=80

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

np :)